OpenStudy (anonymous):

Solve 2x^2 - 9x + 7 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used. Solve 3x^2 - 12x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work.

OpenStudy (anonymous):

Pleasssseee help

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

@mathmale Can you please help me?

OpenStudy (anonymous):

Can you factor an equation when the right side isn't equal to 0? Or are you not allowed to

OpenStudy (anonymous):

I don't know. All I have is the information I already posted

OpenStudy (mathmale):

Leave both equations as they are. You have the following methods to choose from: 1) quadratic formula 2) factoring 3) factoring by grouping

OpenStudy (anonymous):

You could try completing the square

OpenStudy (mathmale):

In both cases you have 0 on the right side, so you can jump right in and attempt factoring. 4) completing the square is another very appropriate method. Thanks, Waffle.

OpenStudy (mathmale):

@BassCatcher15 : Let's catch bass. Choose one of these four methods and show whatever steps of the process you can. Ask questions as appropriate.

OpenStudy (mathmale):

@BassCatcher15 : Please factor x^2-2x+1, for practice.

OpenStudy (anonymous):

I'm clueless right now

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Angel_Halo23

OpenStudy (anonymous):

First Question: Use the cross factoring method, and get (2x-7)(x-1). Set 2x-7=0 and x-1=0. Solve each equation , getting you 7/2 and 1. So x=1, 7/2

OpenStudy (anonymous):

yup negative one

OpenStudy (anonymous):

Negative one? Huh?

OpenStudy (anonymous):

yea i guess

OpenStudy (anonymous):

Okay. So what about the second question?

OpenStudy (anonymous):

hold on a min

OpenStudy (anonymous):

Okayy. :)

OpenStudy (anonymous):

:D hes here

OpenStudy (anonymous):

Great!

OpenStudy (lunymoony):

What do you mean when you use ^ ? times? @BassCatcher15

OpenStudy (lunymoony):

@BassCatcher15

OpenStudy (anonymous):

exponents

OpenStudy (anonymous):

Yeah exponents. OS doesn't allow you to type the small exponent next to the base number, so we have to use the ^ symbol to show an exponent

OpenStudy (mathmale):

Actually, OpenStudy DOES allow you to type in fancy math symbolism:\[2x^2 - 9x + 7 = 0\]

OpenStudy (mathmale):